# PAC Privacy Preserving Diffusion Models

### v1

@@ Abstract
Data privacy protection is garnering increased attention among researchers. Diffusion models (DMs), particularly with strict differential privacy, can potentially produce images with both high privacy and visual quality. However, challenges arise in ensuring robust protection in privatizing specific data attributes, areas where current models often fall short. To address these challenges, we introduce the PAC Privacy Preserving Diffusion Model, a model leverages diffusion principles and ensure Probably Approximately Correct (PAC) privacy. We enhance privacy protection by integrating a private classifier guidance into the Langevin Sampling Process. Additionally, recognizing the gap in measuring the privacy of models, we have developed a novel metric to gauge privacy levels. Our model, assessed with this new metric and supported by Gaussian matrix computations for the PAC bound, has shown superior performance in privacy protection over existing leading private generative models according to benchmark tests.
Machine Learning, ICML
@@ 1 Introduction
Modern deep learning models, fortified with differential privacy as defined by (Dwork et al., 2006), have been instrumental in significantly preserving the privacy of sensitive data (Dwork et al., 2014). DP-SGD (Abadi et al., 2016), a pioneering method for training deep neural networks within the differential privacy framework, applies gradient clipping at each step of the SGD (stochastic gradient descent) to enhance privacy protection effectively.
Diffusion models (DMs) (Song & Ermon, 2019, 2020; Dhariwal & Nichol, 2021) have emerged as state-of-the-art generative models, setting new benchmarks in various applications, particularly in generating high-quality images. When trained under strict differential privacy protocols, these DMs can produce images that safeguard privacy while maintaining high visual fidelity. For instance, DPGEN (Chen et al., 2022) leverages a randomized response technique to privatize the recovery direction in the Langevin MCMC process for image generation. The images produced by DPGEN are not only visually appealing but also compliant with differential privacy standards, although its privacy mechanism has been shown to be data-dependent later on. Moreover, the Differentially Private Diffusion Models (DPDM) (Dockhorn et al., 2022) adapt DP-SGD and introduce noise multiplicity in the training process of diffusion models, demonstrating that DPDM can indeed produce high-utility images while strictly adhering to the principles of differential privacy.
@@
Recently, (Xiao & Devadas, 2023) introduces a novel definition of privacy known as Probably Approximately Correct (PAC) Privacy, representing a significant evolution in privacy-preserving methodologies. This definition diverges from traditional Differential Privacy (DP) by focusing on the difficulty of reconstructing data using any measure function, thereby offering broader applicability beyond the adversarial worst-case scenarios considered by DP. PAC privacy’s strength lies in its model-agnostic nature. It does not depend on an in-depth understanding of a model’s internal workings or its training process. This attribute makes it exceptionally versatile, enabling its application across a wide array of models and contexts with relative ease. Furthermore, leveraging PAC privacy comes with a framework that autonomously determines the minimal noise addition necessary for effective data protection (Xiao & Devadas, 2023). This framework simplifies the process of achieving privacy-preserving data handling, making it more accessible for widespread use in the field of data science and beyond.
@@
  • Through extensive evaluations, we demonstrate that our model sets a new standard in privacy protection, achieving state-of-the-art (SOTA) results, while maintaining image quality at a level that is comparable to other SOTA models.
2 Related Work and Preliminaries
2.1 Differential Privacy(Dwork et al., 2014)
Definition 2.1.
A randomized mechanism $\mathcal{M}$ is said $(\varepsilon,\delta)$- differentially private if for any two adjacent datasets $D$ and $D^{\prime}$ differing in a single datapoint for any subset of outputs $S$:
$\textbf{Pr}[\mathcal{M}(D)\in S]\leq e^{\varepsilon}\cdot\textbf{Pr}[\mathcal{M}(D^{\prime})\in S]+\delta$
@@ 2 Related Work and Preliminaries
2.2 Conditional Image Generation
Dhariwal & Nichol (2021) propose a diffusion model that is enhanced by classifier guidance; it has been shown to outperform existing generative models. By using true labels of datasets, it is possible to train a classifier on noisy images $x_{t}$ at various timesteps $p_{\phi}(y|x_{t},t)$ , and then use this classifier to guide the reverse sampling process $\nabla_{x_{t}}\log p_{\phi}(y|x_{t},t)$. What begins as an unconditional reverse noising process is thus transformed into a conditional one, where the generation is directed to produce specific outcomes based on the given labels
@@ 2.2 Conditional Image Generation
2.3 PAC Privacy (Xiao & Devadas, 2023)
Definition 2.2.
$(\delta,\rho,D)$ PAC Privacy. For a data processing mechanism $\mathcal{M}$, given some data distribution $D$, and a measure function $\rho(.,.)$, we say $\mathcal{M}$ satisfies $(\delta,\rho,D)$-PAC Privacy if the following experiment is impossible: A user generates data $X$ from distribution $D$ and sends $\mathcal{M(X)}$ to an adversary. The adversary who knows $D$ and $\mathcal{M}$ is asked to return an estimation $\tilde{X}\in X^{*}$ on X such that with probability at least $(1-\delta)$, $\rho(\tilde{X},X)=1$.
Definition 2.2 depicts the reconstruction hardness for the attackers to recover the private data distribution $\mathcal{M(X)}$. With a lower bound probability $(1-\delta)$, such that the measure function $\rho(.,.)$ is not able to distinguish the recovery data from the original data. However, the limitation of the naive definition is that the prior distribution of the public dataset is unknown, resulting to the failure of adversarial inferences.
@@ 2.3 PAC Privacy (Xiao & Devadas, 2023)
where $\mathcal{D}_{KL}$ denotes the KL divergence. Utilizing the mutual information, the Definition 2.2 can be rewritten as
Theorem 2.3.
$\mathcal{M}$ satisfies $(\Delta_{KL}\delta,\rho,D)$ PAC Advantage Privacy where
@@
Therefore, the goal for the rest of this paper, following Theorem 2.3, is to ascertain or establish limits for the mutual information $MI(X;M(X))$ with a high level of certainty.
3 Method
@@
$Pr[H(\tilde{x_{i}})=\omega]=\left\{\begin{aligned} &\frac{e^{\epsilon}}{e^{\epsilon}+k-1},&&\omega=x_{i}\\ &\frac{1}{e^{\epsilon}+k-1},&&\omega=x_{i}^{\prime}\in X\setminus x_{i}\end{aligned}\right.$
This mechanism is designed to enforce $\epsilon$-differential privacy by privatizing the recovery direction $d=(\tilde{x}-x)/\sigma^{2}$. Following perturbation, DPGEN learns an energy score function:
$l(\theta,\sigma)=\frac{1}{2}\mathbb{E}_{p(x)}\mathbb{E}_{\tilde{x}\sim\mathcal{N}(x,\sigma^{2})}\left[\left\|\frac{\tilde{x}-x}{\sigma^{2}}-\nabla_{x}\log q_{\theta}(\tilde{x})\right\|^{2}\right]$ (6)
@@
While DPGEN asserts compliance with stringent $\epsilon$-differential privacy, recent findings (Dockhorn et al., 2022) indicate that DPGEN is, in fact, data-dependent. The RR mechanism $M(d)$ in DPGEN holds validity only within the perturbed dataset. Specifically, if an element $z$ belongs to the output set $O$ but not to the perturbed dataset $d$, then $Pr[M(d)=O]=0$, which contravenes the differential privacy definition.
@@ 3.1 Conditional Private Langevin Sampling
$k\Sigma_{\theta}(x_{t-1})\nabla_{x_{t}}logc_{\theta}(y_{i}|x_{t-1})$ (7)
where $k$ is the gradient scale that can be tuned according to the performance of the model, $\Sigma_{\theta}(x_{t-1})$ is the covariance matrix from reverse process Eqn. 2. In Eqn. 7, the term $\nabla_{x_{t}}\log c_{\theta}(y_{i}|x_{t-1})$ directs the Langevin sampling process toward a specific class label $y_{i}$, which is sampled prior to the inference. This private guidance during the inference phase ensures that the synthesized images are protected from privacy breaches related to designated attributes. As a result, the images generated under this guidance have a higher privacy extent when compared to those produced by the original DPGEN.
@@
  Input: class labels $y=\{y_{1},y_{2},...y_{n}\}$ from one of balanced dataset attributes; gradient scale k; $\{\delta_{i}\}_{i=1}^{L}$, $\epsilon$, $T$; pretrained scorenet model $s_{\theta}$;pretrained classifier model $c_{\theta}$
  Randomly sample $y_{i}$ from $y$
  Initialize $x_{0}$
  for $i$ from $1$ to $L$ do
     stepsize $\alpha_{i}$leftarrow $\delta_{i}^{2}/\delta_{L}^{2}$
     for $t$ from $1$ to $T$ do
        Sample noise $z_{t}\sim\mathcal{N}(0,I)$
        $x_{t}$leftarrow $x_{t-1}+\frac{\alpha_{i}}{2}s_{\theta}(x_{t-1},\delta_{i})+\sqrt{\alpha_{i}}z_{t}+k\Sigma_{\theta}(x_{t-1})\nabla_{x_{t}}logc_{\theta}(y_{i}|x_{t-1})$
     end for
     $x_{0}$leftarrow $x_{T}$
  end for
  Return $x_{T}$
@@ 3.2 Privacy Metrics
Most privacy-preserving generative models prioritize assessing the utility of images for downstream tasks, yet often overlook the crucial metric of the models’ own privacy. To bridge this gap in evaluating privacy extent, we have developed a novel algorithm that computes a privacy score for the models obtained. As per Algorithm 2, we commence by preparing images $x_{i}$ generated from Algorithm 1, alongside the original dataset images $x_{k}^{G}$. Subsequently, we process all synthesized images through InceptionV3 (Szegedy et al., 2016).
@@
  Input:images $x_{i}$ generated by algorithm 1; ground truth images $x_{k}^{G}$; sample number $n$; pretrained InceptionV3 model $I_{\theta}$; pretrained classifier model $c_{\theta}$.
  private score $s$leftarrow 0
  for $x_{i}$ from $x_{0}$ to $x_{n}$ do
     find $argmin_{k}||I_{\theta}(x_{i}),I_{\theta}(x_{k}^{G})||_{2}$
     if $c_{\theta}(x_{i})$neq $c_{\theta}(x_{k}^{G})$ then
        $s$leftarrow $s+1$
     end if
  end for
  Return $s/n$
Figure 1: Generated images (the second row) and their closest images from CelebA dataset (the first row).
@@ 3.3 PAC Gaussian Noise Addition Calculation
When the mutual information $MI(X;M(X))$ is insufficient to ensure PAC privacy, additional Gaussian noise $B\sim\mathcal{N}(0,\Sigma_{B})$ can be introduced to yield a reduced mutual information $MI(X;M(X)+B)$ (Xiao & Devadas, 2023). PAC privacy provides a $1-\gamma$ confidence in the noise determination for the deterministic mechanism $M$ to compute the Gaussian noise covariance $\Sigma_{B}$, ensuring that $MI(X;M(X)+B)\leq\nu+\beta$. In practical terms, this means we can evaluate the privacy extent of the model by comparing the mean norm $\mathbb{E}||B||_{2}$ under the same confidence level. A smaller $\mathbb{E}||B||_{2}$ implies a smaller covariance matrix and lesser noise addition, signifying a higher privacy extent of the model.

### v3

# algorithmic

@@ Abstract
Data privacy protection is garnering increased attention among researchers. Diffusion models (DMs), particularly with strict differential privacy, can potentially produce images with both high privacy and visual quality. However, challenges arise such as in ensuring robust protection in privatizing specific data attributes, areas where current models often fall short. To address these challenges, we introduce the PAC Privacy Preserving Diffusion Model, a model leverages diffusion principles and ensure Probably Approximately Correct (PAC) privacy. We enhance privacy protection by integrating a private classifier guidance into the Langevin Sampling Process. Additionally, recognizing the gap in measuring the privacy of models, we have developed a novel metric to gauge privacy levels. Our model, assessed with this new metric and supported by Gaussian matrix computations for the PAC bound, has shown superior performance in privacy protection over existing leading private generative models according to benchmark tests.
@@ 1 Introduction
Modern deep learning models, fortified with differential privacy as defined by Dwork et al. (2006), have been instrumental in significantly preserving the privacy of sensitive data (Dwork et al., 2014). DP-SGD Abadi et al. (2016), a pioneering method for training deep neural networks within the differential privacy framework, applies gradient clipping at each step of the SGD (stochastic gradient descent) to enhance privacy protection effectively.
Diffusion models (DMs) (Song & Ermon, 2019; 2020; Dhariwal & Nichol, 2021) have emerged as state-of-the-art generative models, setting new benchmarks in various applications, particularly in generating high-quality images. When trained under strict differential privacy protocols, these DMs can produce images that safeguard privacy while maintaining high visual fidelity. For instance, DPGEN (Chen et al., 2022) leverages a randomized response technique to privatize the recovery direction in the Langevin MCMC process for image generation. The images produced by DPGEN are not only visually appealing but also compliant with differential privacy standards, although its privacy mechanism has been shown to be data-dependent later on. Moreover, the Differentially Private Diffusion Models (DPDM) (Dockhorn et al., 2022) adapt DP-SGD and introduce noise multiplicity in the training process of diffusion models, demonstrating that DPDM can indeed produce high-utility images while strictly adhering to the principles of differential privacy.
@@
Recently, Xiao & Devadas (2023) introduces a novel definition of privacy known as Probably Approximately Correct (PAC) Privacy, representing a significant evolution in privacy-preserving methodologies. This definition diverges from traditional Differential Privacy (DP) by focusing on the difficulty of reconstructing data using any measure function, thereby offering broader applicability beyond the adversarial worst-case scenarios considered by DP. PAC privacy’s strength lies in its model-agnostic nature: it does not depend on an in-depth understanding of a model’s internal workings or its training process. This attribute makes it exceptionally versatile, enabling its application across a wide array of models and contexts with relative ease. Furthermore, leveraging PAC privacy comes with a framework that autonomously determines the minimal noise addition necessary for effective data protection (Xiao & Devadas, 2023). This framework simplifies the process of achieving privacy-preserving data handling, making it more accessible for widespread use in the field of data science and beyond.
@@
  • Through extensive evaluations, we demonstrate that our model sets a new standard in privacy protection of specific attributes, achieving state-of-the-art (SOTA) results, while maintaining image quality at a level that is comparable to other SOTA models.
2 Related Work
2.1 Early Works on Differentially Private Image Generation
Image Synthesis with differential privacy has been studied extensively during the research. Recently, there has been an increased emphasis on utilizing sophisticated generative models to improve the quality of differentially private synthetic data (Hu et al., 2023). Some approaches employ Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), or GANs that have been trained using the Private Aggregation of Teacher Ensembles (PATE) framework (Xie et al., 2018; Chen et al., 2020; Harder et al., 2021; Torkzadehmahani et al., 2019). Other contributions leverage variational autoencoders(VAEs) (Pfitzner & Arnrich, 2022; Jiang et al., 2022; Takagi et al., 2021), or take advantage of customized architectures (Cao et al., 2021; Harder et al., 2023). However, there are several limitations for those DP synthesizers: (1) Failure when applying to high-dimensional data, primarily due to the constraints imposed by discretization. (2) Limited image quality and lack of expressive generator networks (Cao et al., 2021).
2.2 Differentially Private Diffusion Models
Diffusion models (DMs) (Song & Ermon, 2019; 2020; Dhariwal & Nichol, 2021) , recognized for setting new standards in image generation, can produce high-quality, privacy-compliant images when trained with differential privacy protocols. For example, DPGEN (Chen et al., 2022) employs a data-dependant randomized response method to privatize the recovery direction in the Langevin MCMC process for image generation. Furthermore, Differentially Private Diffusion Models (DPDM) (Dockhorn et al., 2022), which adapt DP-SGD and introduce noise multiplicity, both demonstrate the feasibility of generating visually appealing, privacy-protective images. Subsequent advancements, including fine-tuning existing models and employing novel diffusion model architectures, have been made to boost the effectiveness of differentially private image generation, as detailed in Ghalebikesabi et al. (2023); Lyu et al. (2023). Nevertheless, as previously noted in the introduction, there remains three key challenges to be addressed. In the following sections, we propose solutions and methods to tackle these issues.
3 Background
3.1 Differential Privacy
A randomized mechanism $\mathcal{M}$ is said $(\varepsilon,\delta)$- differentially private if for any two adjacent datasets $D$ and $D^{\prime}$ differing in a single datapoint for any subset of outputs $S$ as follows (Dwork et al., 2014):
$\textbf{Pr}[\mathcal{M}(D)\in S]\leq e^{\varepsilon}\cdot\textbf{Pr}[\mathcal{M}(D^{\prime})\in S]+\delta$ (1)
@@ 3.1 Differential Privacy
3.2 Conditional Diffusion Models
Dhariwal & Nichol (2021) proposed a diffusion model that is enhanced by classifier guidance; it has been shown to outperform existing generative models. By using true labels of datasets, it is possible to train a classifier on noisy images $x_{t}$ at various timesteps $p_{\phi}(y|x_{t},t)$ , and then use this classifier to guide the reverse sampling process $\nabla_{x_{t}}\log p_{\phi}(y|x_{t},t)$. What begins as an unconditional reverse noising process is thus transformed into a conditional one, where the generation is directed to produce specific outcomes based on the given labels
@@ 3.2 Conditional Diffusion Models
3.3 PAC Privacy
Definition 3.3.1. ($(\delta,\rho,D)$ PAC Privacy (Xiao & Devadas, 2023)). For a data processing mechanism $\mathcal{M}$, given some data distribution $D$, a measure function $\rho(.,.)$, and a finite set $X^{*}$, we say $\mathcal{M}$ satisfies $(\delta,\rho,D)$-PAC Privacy if the following experiment is impossible: A user generates data $X$ from distribution $D$ and sends $\mathcal{M(X)}$ to an adversary. The adversary who knows $D$ and $\mathcal{M}$ is asked to return an estimation $\tilde{X}\in X^{*}$ on X such that with probability at least $(1-\delta)$, $\rho(\tilde{X},X)=1$.
Definition above depicts the reconstruction hardness for the attackers to recover the private data distribution $\mathcal{M(X)}$. With a lower bound probability $(1-\delta)$, such that the measure function $\rho(.,.)$ is not able to distinguish the recovery data from the original data. However, the limitation of the naive definition is that the prior distribution of the public dataset is unknown, resulting to the failure of adversarial inferences.
@@ 3.3 PAC Privacy
where $\mathcal{D}_{KL}$ denotes the KL divergence. Utilizing the mutual information, the Definition can be rewritten as $\mathcal{M}$ satisfies $(\Delta_{KL}\delta,\rho,D)$ PAC Advantage Privacy where
@@
Therefore, the goal for the rest of this paper, following Theorem above, is to ascertain or establish limits for the mutual information $MI(X;M(X))$ with a high level of certainty.
4 Methods
@@
$Pr[H(\tilde{x_{i}})=x_{i}^{r}]=\left\{\begin{aligned} &\frac{e^{\epsilon}}{e^{\epsilon}+k-1},&&\omega=x_{i}\\ &\frac{1}{e^{\epsilon}+k-1},&&\omega=x_{i}^{\prime}\in X\setminus x_{i}\end{aligned}\right.$
With the method, $\tilde{x_{i}}$ is designed to point to one of its $k$ nearest neighbors with certain probability $Pr[H(\tilde{x_{i}})=x_{i}^{r}]$, which is designed to enforce $\epsilon$-differential privacy by privatizing the recovery direction $d=(\tilde{x_{i}}-x_{i}^{r})/\sigma^{2}_{i}$. Following perturbation, DPGEN learns an energy score function:
$l(\theta,\sigma)=\frac{1}{2}\mathbb{E}_{p(x)}\mathbb{E}_{\tilde{x}\sim\mathcal{N}(x,\sigma^{2})}\left[\left\|d-\nabla_{x}\log q_{\theta}(\tilde{x})\right\|^{2}\right]$ (7)
@@
While DPGEN asserts compliance with stringent $\epsilon$-differential privacy, recent findings Dockhorn et al. (2022) indicate that DPGEN is, in fact, data-dependent. The RR mechanism $M(d)$ in DPGEN holds validity only within the perturbed dataset. Specifically, if an element $z$ belongs to the output set $O$ but not to the perturbed dataset $d$, then $Pr[M(d)=O]=0$, which contravenes the differential privacy definition.
@@ 4.1 Conditional Private Langevin Sampling
$k\Sigma_{\theta}(x_{t-1})\nabla_{x_{t-1}}\log c_{\theta}(y_{n}|x_{t-1})$ (8)
where $k$ is the gradient scale that can be tuned according to the performance of the model, $\Sigma_{\theta}(x_{t-1})$ is the covariance matrix from reverse process Eqn. 3. In Eqn. 8, the term $\nabla_{x_{t-1}}\log c_{\theta}(y_{n}|x_{t-1})$ directs the Langevin sampling process toward a specific class label $y_{n}$, which is sampled prior to the inference. This private guidance during the inference phase ensures that the synthesized images are protected from privacy breaches related to designated attributes.
In essence, private classifier guidance can be interpreted as a method of deliberately altering certain generated trajectories. To illustrate, consider an image that is autonomously generated to depict Celebrity A donning eyewear, thereby reflecting the known fact that Celebrity A habitually wears glasses. However, when sampling is conducted with the application of private classifier guidance, the attributes are subject to variation. Consequently, there is a possibility that Celebrity A may be depicted without eyewear in the resulting image. This alteration effectively safeguards the attribute of ’glasses’ from being a consistent feature in the generated representation of Celebrity A. As a result, the images generated under this guidance have a higher privacy extent when compared to those produced by the original DPGEN.
@@
0:  class labels $y=\{y_{1},y_{2},...y_{n}\}$ from one of balanced dataset attributes; gradient scale k; $\{\delta_{i}\}_{i=1}^{L}$, $\epsilon$, $T$; pretrained scorenet model $s_{\theta}$;pretrained classifier model on noisy images $c_{\theta}$
1:  Randomly sample $y_{n}$ from $y$
2:  Initialize $x_{0}$
3:  for $i$ from $1$ to $L$ do
4:     stepsize $\alpha_{i}$leftarrow $\delta_{i}^{2}/\delta_{L}^{2}$
5:     for $t$ from $1$ to $T$ do
6:        Sample noise $z_{t}\sim\mathcal{N}(0,I)$
7:        $x_{t}$leftarrow $x_{t-1}+\frac{\alpha_{i}}{2}s_{\theta}(x_{t-1},\delta_{i})+\sqrt{\alpha_{i}}z_{t}+k\Sigma_{\theta}(x_{t-1})\nabla_{x_{t-1}}\log c_{\theta}(y_{n}|x_{t-1})$
8:     end for
9:     $x_{0}$leftarrow $x_{T}$
10:  end for
11:  Return $x_{T}$
@@ 4.2 Privacy Metrics
Most privacy-preserving generative models prioritize assessing the utility of images for downstream tasks, yet often overlook the crucial metric of the models’ own privacy. To bridge this gap in evaluating privacy extent, we have developed a novel algorithm that computes a privacy score for the models obtained. As per Algorithm 2, we commence by preparing images $x_{i}$ generated from Algorithm 1, the original dataset images $x_{k}^{G}$, and alongside the classifier model which trained seperately with clean images (different from the classifier from Algorithm 1). Subsequently, we process all synthesized images through InceptionV3 (Szegedy et al., 2016).
@@
0:  images $x_{i}$ generated by algorithm 1; ground truth images $x_{k}^{G}$; sample number $n$; pretrained InceptionV3 model $I_{\theta}$; pretrained classifier model $c_{\theta}$.
1:  private score $s$leftarrow 0
2:  for $x_{i}$ from $x_{0}$ to $x_{n}$ do
3:     find $argmin_{k}||I_{\theta}(x_{i}),I_{\theta}(x_{k}^{G})||_{2}$
4:     if $c_{\theta}(x_{i})$neq $c_{\theta}(x_{k}^{G})$ then
5:        $s$leftarrow $s+1$
6:     end if
7:  end for
8:  Return $s/n$
@@ 4.3 PAC Gaussian Noise Addition Calculation
When the mutual information $MI(X;M(X))$ is insufficient to ensure PAC privacy, additional Gaussian noise $B\sim\mathcal{N}(0,\Sigma_{B})$ can be introduced to yield a reduced mutual information $MI(X;M(X)+B)$ (Xiao & Devadas, 2023). PAC privacy provides a $1-\gamma$ confidence in the noise determination for the deterministic mechanism $M$ shown in Algorithm 3, to compute the Gaussian noise covariance $\Sigma_{B}$, ensuring that $MI(X;M(X)+B)\leq\nu+\beta$. In practical terms, this means we can evaluate the privacy extent of the model by comparing the mean norm $\mathbb{E}||B||_{2}$ under the same confidence level. A smaller $\mathbb{E}||B||_{2}$ implies a smaller covariance matrix and lesser noise addition, signifying a higher privacy extent of the model.
Algorithm 3 (1 - $\gamma$)-Confidence Noise Determination of Deterministic Mechanism
0:  Diffusion-Privacy model $M$ , data distribution $D$, sampling complexity $m$, security parameter $c$, and mutual information quantities $\nu$ and $\beta$.
1:  Train $M$ model with data $X$.
2:  for $k=1,2,\ldots,m$ do
3:     sample images $y^{(k)}=M(X^{(k)})$.
4:  end for
5:  Calculate empirical mean $\hat{\mu}=\frac{1}{m}\sum_{k=1}^{m}y^{(k)}$ and the empirical covariance estimation $\hat{\Sigma}=\frac{1}{m}\sum_{k=1}^{m}(y^{(k)}-\hat{\mu})(y^{(k)}-\hat{\mu})^{T}$.
6:  Apply singular value decomposition (SVD) on $\hat{\Sigma}$ and obtain the decomposition as $\hat{\Sigma}=\hat{U}\hat{\Lambda}\hat{U}^{T}$, where $\hat{\Lambda}$ is the diagonal matrix of eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{d}$.
7:  Determine the maximal index $j_{0}=\arg\max_{j}\lambda_{j}$ for those $\lambda_{j}>c$.
8:  if $\min_{1\leq j\leq j_{0},1\leq d}(\lambda_{j}-\hat{\lambda}_{j})>r\sqrt{d/c+2c}$ then then
9:     for $j=1,2,\ldots,d$ do
10:        Determine the $j$-th element of a diagonal matrix $A_{B}$ as
$\lambda_{B,j}=\frac{2\nu}{\sqrt{\lambda_{j}+10c\nu/\beta}\cdot\left(\sum_{j=1}^{d}\sqrt{\lambda_{j}+10c\nu/\beta}\right)}$
11:     end for
12:     Determine the Gaussian noise covariance as $\Sigma_{B}=\hat{U}A_{B}\hat{U}^{T}$.
13:  else
14:     Determine the Gaussian noise covariance as $\Sigma_{B}=\left(\sum_{j=1}^{d}\lambda_{j}+dc/(2\nu)\right)\cdot I_{d}$.
15:  end if
16:  Return Gaussian covariance matrix $\Sigma_{B}$.